A die is rolled twice
What is the probabilty of both being a 6

Answer by stanbon(75887)

The results of the two rolls are independent; that means what happens
on the 1st roll has no effect on what happens on the 2nd roll.
So Prob( 6 and then 6) = Prob 6 * Prob 6 = 1/6 *1/6 = 1/36
------------
Cheers,
Stan H.